OpenStudy (anonymous):

(6.3*10^11)/(9*10^5) show the expression in simplified form

OpenStudy (anonymous):

In this case, simplified form would just be getting the answer (remember to use order of operations)-- Lets look at the first half of the equation first. (6.3*10^11) 10^11=100,000,000,000 100,000,000,000*6.3=630,000,000,000 this leaves you with (630,000,000,000)/(9*10^5) as your equation then start work on the other side 10^5=100,000 100,000*9=900,000 your equation now is (630,000,000,000)/(900,000) divide it and your answer is 700,000